A body is projected downwards at an angle 30degrees to the horizontal with a velocity 9.8m/s from the top of a 29.4 m high tower. Its horizontal distance on the ground from the tower is 1)9.8 2)9.3√3 3)4.9

Grade:12

1 Answers

Arun
6 years ago
initial vertical velocity= usin30 = 4.9m/s
vertical distance traversed= 29.4m
So,
h = ut + 1/2gt2
29.4=4.9*t +4.9t2
on solving we get
t2 + t – 6 = 0 (on solving quadratic equation)
t = 3 seconds.
Now you can also solve for distance.
 
Regards
Arun (askIITians forum expert)